### v1

# Similarity-Distance-Magnitude Activations

Abstract
We introduce a more robust and interpretable formulation of the standard softmax activation function commonly used with neural networks by adding Similarity (i.e., correctly predicted depth-matches into training) awareness and Distance-to-training-distribution awareness to the existing output Magnitude (i.e., decision-boundary) awareness. When used as the final-layer activation with language models, the resulting Similarity-Distance-Magnitude (sdm) activation function is more robust than the softmax function to co-variate shifts and out-of-distribution inputs in high-probability regions, and provides interpretability-by-exemplar via dense matching. Complementing the prediction-conditional estimates, the sdm activation enables a partitioning of the class-wise empirical CDFs to guard against low class-wise recall among selective classifications. These properties make it preferable for selective classification, even when considering post-hoc calibration methods over the softmax.
1 Introduction
Large neural-network-based language models (LLMs) pose a challenge for interpretable and reliable deployment given the non-identifiability of their parameters (Hwang & Ding, 1997, inter alia) [^1] Informally, this means that two or more distinct sets of values for the parameters can result in identical output distributions. As a consequence, interpreting the parameters of such models is typically much more complicated than with a simple linear regression model, for example., which can number in the billions or more. Instead of directly interpreting parameters, we might instead consider moving the focus of interpretation to auditing predictions as a form of interpretability by example, or exemplar, over the representation space of such models via dense matching. However, for real-world deployments, robust approaches for predictive uncertainty are also needed, both for human decision-making and for constructing sequentially dependent LLM pipelines.
Known theoretical results limit the statistical quantities that can be derived over LLMs. Statistical assurances in the distribution-free setting are limited to approximately conditional quantities (Valiant, 1984; Lei & Wasserman, 2014; Foygel Barber et al., 2020, inter alia). Further, even typical approximately conditional quantities can be difficult to obtain in practice, since the minimal assumption of exchangeability with a known held-out data set is itself often violated with co-variate and label shifts, which can be difficult to foresee with existing methods. Epistemologically, the prevalence of hallucinations and highly-confident wrong answers with widely deployed LLMs suggests a technical impasse in effectively modeling the predictive uncertainty, despite significant work from Bayesian, Frequentist, and empirically motivated perspectives (Gal & Ghahramani, 2016; Angelopoulos et al., 2021; Guo et al., 2017; Lakshminarayanan et al., 2017; Ovadia et al., 2019, inter alia). A foundational piece is evidently missing from the picture.
Given these intrinsic challenges, we approach the problem of uncertainty quantification over LLMs from a new angle and ask: Can we leverage the metric learning and dense matching capabilities of neural networks over high-dimensional inputs to at least aim to maximize, with minimal distributional assumptions, the separation of aleatoric (irreducible) uncertainty and epistemic (reducible) uncertainty, decomposing the sources of the latter in a manner that is interpretable and actionable?
We answer this question in the affirmative with a conceptually parsimonious, LLM-driven partitioning of the data to decompose sources of epistemic uncertainty: Correctly predicted depth-matches into the training set ($\operatorname{\text{{Similarity}}}$), the $\operatorname{\text{{Distance}}}$ to the training set, and the distance to the decision-boundary ($\operatorname{\text{{Magnitude}}}$). We use these signals to construct a new activation function, the $\operatorname{\text{{sdm}}}$ activation, which can be used as a replacement for the standard $\operatorname{softmax}$ operation, as, for example, the final-layer activation. The $\operatorname{\text{{sdm}}}$ activation enables more reliable estimates of the predictive uncertainty for selective classification (Chow, 1957; Geifman & El-Yaniv, 2017, inter alia), which addresses the need for uncertainty quantification with LLMs as part of multi-stage decision pipelines, in settings subject to co-variate shifts and out-of-distribution inputs.
In summary, in this work:
  • We introduce the $\operatorname{\text{{Similarity}}}$-$\operatorname{\text{{Distance}}}$-$\operatorname{\text{{Magnitude}}}$ ($\operatorname{\text{{sdm}}}$) activation function, which encodes strong signals of epistemic uncertainty, to replace the $\operatorname{softmax}$ operation.
  • We specifically examine the behavior of the $\operatorname{\text{{sdm}}}$ activation as a final-layer activation over language models for selective classification. Even when considering typical post-hoc calibration methods, the $\operatorname{\text{{sdm}}}$ activation is more robust than the $\operatorname{softmax}$ activation to co-variate-shifted and out-of-distribution inputs, ceteris paribus, while remaining informative over in-distribution data.
  • We introduce an approach for controlling the recall of selective classifications, based on a data-driven partitioning of the class-wise CDFs via the $\operatorname{\text{{sdm}}}$ activation. Combined with the empirically-robust prediction-conditional estimates from the $\operatorname{\text{{sdm}}}$ activation, this results in a straightforward, and empirically achievable, quantity of interest for selective classification.
2 Preliminaries
2.1 Setting
We consider the standard multi-class classification setting (e.g., predicting the binary sentiment of a movie review). We are given a training dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}=\{({\bm{x}}_{n},y_{n})\}_{n=1}^{N}$ of inputs, ${\bm{x}}\in{\mathcal{X}}$, paired with their corresponding ground-truth discrete labels, $y\in{\mathcal{Y}}=\{1,\ldots,C\}$, and a labeled calibration dataset, $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, drawn from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. We are then given a new test instance, $\bm{x}$}, from an unlabeled test set, $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, and seek to estimate the label with a prediction, $\hat{y}$, via the un-normalized log probabilities (“logits”, informally) of a final linear layer: ${\bm{z}}={\bm{W}}^{T}{\bm{h}}+{\bm{b}}$, where ${\bm{h}}=\operatorname{{\rm{network}}}({\bm{x}};\theta)$ is the final hidden state of a network parameterized by $\theta$. The discrete prediction is taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}}$; however, for learning $\theta$, $\bm{W}$}, and $\bm{b}$}, and for human decision-making, we also seek an estimate of the predictive uncertainty, $p(y\,|\,{\bm{x}})$, which is typically obtained by normalizing $\bm{z}$} via the $\operatorname{softmax}$ operation described next. We will make a distinction between models, $\mathcal{M}$} (defined by $\theta$, $\bm{W}$}, and $\bm{b}$}, and when applicable, the exemplar adaptor, described below), which produce the prediction, $\hat{y}$, and estimators, $\mathcal{E}$}, which provide an estimate of $p(y\,|\,{\bm{x}})$, because different estimators can be used over the same model.
2.2 Softmax and the Cross-Entropy loss
The $\operatorname{softmax}$ activation is commonly used in neural network architectures, including, for example, as a router in self-attention mechanisms (Vaswani et al., 2017) and mixture-of-experts models (Shazeer et al., 2017), and forming the basis of the cross-entropy loss used for next-token training of LLMs. It is the typical final output layer of LLMs, converting the un-normalized model logits to a normalized probability distribution:
$\operatorname{softmax}({\bm{z}})_{i}=\frac{e^{\tau\cdot z_{i}}}{\sum^{C}_{c=1}{e^{\tau\cdot z_{c}}}},1\leq i\leq C,\tau\geq 0$ (1)
The inverse-temperature parameter, $\tau$, controls the sharpness of the distribution. As $\tau\rightarrow 0$, the output of $\operatorname{softmax}({\bm{z}})$ converges to a uniform distribution where each class has probability $\frac{1}{C}$; as $\tau\rightarrow\infty$, the output converges to a distribution in which all of the mass is assigned to a single class. In deep learning, $\tau$ is treated as a learnable, global hyper-parameter; instance-wise variation in the distance to the decision-boundary is thus determined by the relative $\operatorname{\text{{Magnitude}}}$ of $z_{\hat{y}}$. This model is learned by minimizing the cross-entropy loss between $\bm{z}$} and the index of the true labels over $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. The natural logarithm of the loss is the counterpart to the base $e$ of the $\operatorname{softmax}$:
${\mathcal{L}}(\theta,{\bm{W}},{\bm{b}};\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}})=-\frac{1}{N}\sum_{n}^{N}\log_{e}\left(\frac{e^{\tau\cdot z_{y_{n}}}}{\sum^{C}_{c=1}{e^{\tau\cdot z_{c}}}}\right)$ (2)
3 Methods
In this work, we revisit Eq. 1 and  2. We seek to decouple the sources of epistemic uncertainty via a new activation function that is conceptually:
$\operatorname{\text{{sdm}}}({\bm{z}})_{i}=\frac{{\operatorname{\text{{Similarity}}}}^{\operatorname{\text{{Distance}}}\cdot\operatorname{\text{{Magnitude}}}_{i}}}{\sum^{C}_{c=1}{{\operatorname{\text{{Similarity}}}}^{\operatorname{\text{{Distance}}}\cdot\operatorname{\text{{Magnitude}}}_{c}}}}$ (3)
with a corresponding negative log likelihood loss that takes into account the change of base (§ 3.1). Unique to this setting, a modification to label-conditional conformal prediction (Vovk et al., 2005) then follows via a parsimonious partitioning of the class-wise empirical CDFs, providing a principled basis for controlling the recall among selective classifications, combined with empirically-robust prediction-conditional estimates.
3.1 Similarity-Distance-Magnitude Activation Functions
Calculating the $\operatorname{\text{{sdm}}}$ activation involves training an exemplar adaptor, a 1-D CNN adaptor (with a final linear layer) over the frozen hidden states of a network, to induce distilled, compressed representations of the underlying network’s representation space conditional on its predictions. The resulting representations provide a probabilistic mapping to the training, or support, set. In this way, neural networks, including large pre-trained networks, can be viewed as hidden instance-based metric learners, from which we can then derive signals of the epistemic uncertainty.
3.1.1 Exemplar Adaptor
We take as the CNN of our exemplar adaptor $g:{\bm{h}}\in\mathbb{R}^{D}\mapsto{\bm{h}}^{\prime}\in\mathbb{R}^{M}$, a 1-D CNN that takes as input $\bm{h}$} of the underlying network. The CNN has $M$ filters, the filter applications of which produce $\bm{h}$$^{\prime}$, the distilled representation of the underlying network. A final linear layer, ${\bm{z}}^{\prime}={\bm{W}}^{\prime T}{\bm{h}}^{\prime}+{\bm{b}}^{\prime},{\bm{z}}^{\prime}\in\mathbb{R}^{C}$, then replaces the underlying network’s linear layer, with the discrete prediction taken as $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}^{\prime}}$. This exemplar adaptor will then enable us to derive the key signals of epistemic uncertainty, $\operatorname{\text{{Similarity}}}$, $\operatorname{\text{{Distance}}}$, and $\operatorname{\text{{Magnitude}}}$ described next.
3.1.2 $\operatorname{\text{{Similarity}}}$
[…]
$q=\sum^{|\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}|}_{i=1}\mathbf{1}_{\hat{y}=\hat{y}^{\rm{tr}}_{(i)}}\cdot\mathbf{1}_{\hat{y}^{\rm{tr}}_{(i)}=y^{\rm{tr}}_{(i)}}\cdot\mathbf{1}_{i-1=\sum^{i-1}_{j=1}\mathbf{1}_{\hat{y}=\hat{y}^{\rm{tr}}_{(j)}}\cdot\mathbf{1}_{\hat{y}^{\rm{tr}}_{(j)}=y^{\rm{tr}}_{(j)}}}$ (4)
where the rightmost indicator function, $\mathbf{1}\in\{0,1\}$, ensures consecutive (depth-wise) matches. [^3] This seemingly simple rule differs from traditional KNN rules (Cover & Hart, 1967; Devroye et al., 1996, inter alia) in two critical respects: The neural network serves as a semi-supervised metric learner of the distances between the dense representations that identify the instances, and there is a model prediction (in addition to the ground-truth label) for each instance in the support set. The former enables effective partitioning despite the curse of high dimensions; the latter provides an additional indicator of reliability for each instance. By definition, $\operatorname{\text{$q$}}$ cannot exceed the count of the most prevalent class label in $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, and since we assume an approximately equal number of points for each class, $\operatorname{\text{$q$}}\ll|\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}|$ is typical. For the special case of calculating $q$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, which only occurs during learning, we exclude the self-match.
3.1.3 $\operatorname{\text{{Distance}}}$
[…]
$\operatorname{\text{$d$}}=\min\left[1-{\rm{eCDF}}^{y_{1}}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}}),\ldots,1-{\rm{eCDF}}^{y_{C}}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})\right]$ (5)
The empirical CDFs are determined by the labeled points in $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for which $\operatorname{\text{$q$}}>0$, where, as indicated by the superscripts, the stratification of points is by the true labels, $y$. For example, ${\rm{eCDF}}^{y_{1}}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})$ is the empirical CDF of $\operatorname{\text{$d_{\rm{nearest}}$}}$ values in $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for which $y=1$, a notation convention we will use throughout. (Points with $\operatorname{\text{$q$}}=0$ are effectively out-of-distribution points and treated as such in downstream decision-making, so they are excluded to avoid biasing the estimates.) At test time, we do not see $y$; instead, the minimum is calculated over the quantiles of each of the class-conditional eCDFs, regardless of $\hat{y}$. As with $q$, for the special case of calculating $\operatorname{\text{$d$}}$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, we replace ${\rm{eCDF}}^{y_{c}}_{\rm{ca}}$ with the analogous ${\rm{eCDF}}^{y_{c}}_{\rm{tr}}$, the class-wise empirical CDFs of $\operatorname{\text{$d_{\rm{nearest}}$}}$ over $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ excluding self-matches.
3.1.4 $\operatorname{\text{{Magnitude}}}$
We take as the $\operatorname{\text{{Magnitude}}}$, or distance to the decision boundary, $z^{\prime}_{\hat{y}}$, as in the standard $\operatorname{softmax}$ case but via $\bm{z}$$^{\prime}$ from the linear layer of the exemplar adaptor.
3.1.5 SDM Activation: Formulation
We use the above quantities to define the $\operatorname{\text{{sdm}}}$ activation function:
$\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{i}=\frac{(2+q)^{\operatorname{\text{$d$}}\cdot z^{\prime}_{i}}}{\sum^{C}_{c=1}{(2+q)^{\operatorname{\text{$d$}}\cdot z^{\prime}_{c}}}},1\leq i\leq C$ (6)
The output distribution becomes sharper with higher values of $\operatorname{\text{$q$}}$, $\operatorname{\text{$d$}}$, and $z^{\prime}$. Also note that when $\operatorname{\text{$d_{\rm{nearest}}$}}$ exceeds the largest distance observed in the labeled data, $\operatorname{\text{$d$}}=0$ and the output distribution is uniform, reflecting a maximally high (i.e., out-of-distribution) epistemic uncertainty estimate. The standard $\operatorname{softmax}$ with $\tau=1$ is recovered by setting q=e2,d=1q=e-2,d=1. As with the $\operatorname{softmax}$ operation, $\operatornamewithlimits{arg\,max}\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})=\operatornamewithlimits{arg\,max}{\bm{z}}^{\prime}$.
3.1.6 SDM Activation: Loss and Training
A loss analogous to Eq. 2 then follows with the applicable change of base. We use this loss to train the weights of the exemplar adaptor, which includes the parameters of the linear layer ($\bm{W}$$^{\prime}$ and $\bm{b}$$^{\prime}$), as well as the convolution weights and biases, which we collectively represent with $\bm{G}$}. The weights of the underlying $\operatorname{{\rm{network}}}$ remain fixed.
${\mathcal{L}}({\bm{G}},{\bm{W}}^{\prime},{\bm{b}}^{\prime};\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}})=-\frac{1}{N}\sum_{n}^{N}\log_{(2+q)}\left(\frac{(2+q)^{\operatorname{\text{$d$}}\cdot z^{\prime}_{y_{n}}}}{\sum^{C}_{c=1}{(2+q)^{\operatorname{\text{$d$}}\cdot z^{\prime}_{c}}}}\right)$ (7)
The first epoch of training is initialized with a standard $\operatorname{softmax}$ (i.e., setting q=e2,d=1q=e-2,d=1). Training then proceeds by re-calculating $q$ and $d$ for each ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ after each epoch. We take as the stopping criteria for one learning round as the epoch with the highest balanced (across classes) average loss over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. We repeat this process for $J$ iterations of random shuffles and splits of $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ and parameter initializations, choosing the final model as that with the globally highest balanced (across classes) average loss over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$.
3.2 Evaluating Selective Classification
As a common, unambiguous baseline quantity for comparing selective classifiers, we seek a straightforward balance between precision and recall. Among the selective classifications from an estimator, we seek (I) prediction-conditional accuracy and (II) class-conditional accuracy above a given threshold, $\alpha\in(\frac{1}{C},1]$.
To evaluate this quantity, we only consider the points for which the given estimator assigns a high-probability of at least $\alpha$, which is typically near 1, such as $\alpha=0.95$ in our experiments. We refer to this set of points as the admitted, or non-rejected, set. Then, given ground-truth values for $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, we assess whether the conditional accuracies of the admitted set are at least $\alpha$ when (I) stratifying by the predicted labels, $\hat{y}$, and when (II) stratifying by the true labels, $y$.
The estimator that rejects all points would meet these conditions. However, given two estimators that meet these conditions, we prefer that which rejects fewer points, ceteris paribus. In other words, we seek estimators that meet our reliability condition and are informative (i.e., maximize the number of points that are properly admitted), but when the estimator is uncertain, we prefer rejection over unexpectedly falling under the desired $\alpha$ probability threshold.
Quantity I corresponds to top-label calibration (Gupta & Ramdas, 2022), but with a single bin for evaluation, $\alpha,1]$, removing ambiguity with regard to the choice of binning the probabilities. Quantity II does not directly correspond to quantities typically examined in the empirical calibration literature (Brier, 1950; Dawid, 1982; Guo et al., 2017; Vaicenavicius et al., 2019; Kull et al., 2019, inter alia), but it approximates444 For simplicity of presentation, we omit consideration of the Beta-distributed error term that is a function of the effective sample size of split-conformal coverage (Vovk, 2012). In practice, this term is negligible in the present experiments given $|\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}|$ and the resolution of the comparisons. label-conditional conformal coverage in the restricted setting of class-wise thresholds that only admit prediction sets of cardinality 1. We describe a straightforward procedure to estimate this quantity next.
3.2.1 Controlling Selective Classification Recall
In general, the statistical coverage assurance of marginal split-conformal estimators does not directly correspond to selection, since the coverage assurance is not conditional on the set size. We may instead seek one of various approximately conditional notions of coverage (Romano et al., 2020; Angelopoulos et al., 2021, inter alia); however, there is no guarantee that when we stratify by sets of cardinality 1, coverage will be maintained. However, there is a restricted setting in which label-conditional conformal estimators do provide a meaningful notion of class-conditional coverage for selection. Assuming $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ are exchangeable, if the conformity score for each label is from a categorical distribution and the resulting thresholding of the class-wise empirical CDFs results in class-wise thresholds that are all greater than $\frac{1}{C}$, then the cardinality 1 sets will, on average, obtain class-conditional coverage, by definition. [^5] Although the formal interpretations are not identical, the evaluation of class-conditional coverage for a single estimate of such cardinality 1 prediction sets in this restricted setting over $\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ is numerically equivalent to assessing the class-conditional accuracy, when not considering the sample size error term. Unfortunately, it may be rare to encounter this restricted setting over the full data distributions of real-world tasks. Instead, we will use the $\operatorname{\text{{sdm}}}$ activation to estimate a partitioning of the distribution into a region that approximately fulfills these assumptions.
First, we rescale the $\operatorname{\text{{Similarity}}}$ estimate to take into account the $\operatorname{\text{{Distance}}}$ and $\operatorname{\text{{Magnitude}}}$, given the predicted class. The resulting value [^6] The $\min$ ensures that $q=0$ instances remain 0. will be the basis for partitioning the distribution:
[…]
Next, we estimate label-conditional conformal thresholds, $\psi_{1},\ldots\psi_{C}]$, over the output from the $\operatorname{\text{{sdm}}}$ activation among a subset of the distribution determined by progressively larger values of $q^{\prime}$ (among $\left\lfloor{q^{\prime}}\right\rfloor>0$) until all thresholds are at least $\alpha$. By setting our stopping criteria at $\alpha$ rather than $\frac{1}{C}$, we also restrict the region to our empirically-motivated prediction-conditional quantity, $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{\hat{y}}$. The procedure appears in Alg. 1. If we find a finite $q^{\prime}_{\rm{min}}$ that obtains such thresholds, we refer to the resulting region as the $\operatorname{\text{{high-reliability}}}$ ($\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$) region, taking membership in this region as our selection criteria:
$\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}:=\begin{cases}\hat{y}&\text{if~}q^{\prime}\geq q^{\prime}_{\rm{min}}\wedge\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{\hat{y}}\geq\psi_{\hat{y}}\\ \bot&\text{otherwise}\end{cases}$ (9)
where $\bot$ indicates a rejected (non-admitted) point and $\hat{y}=\operatornamewithlimits{arg\,max}{{\bm{z}}^{\prime}}$.
To calculate this quantity for new, unseen test points ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$, we require $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ to calculate $\operatorname{\text{$q$}}$ and $\operatorname{\text{$d_{\rm{nearest}}$}}$; the cached class-wise empirical CDFs of the distances over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ of Eq. 5; and $q^{\prime}_{\rm{min}}$ and the thresholds ($\psi_{1},\ldots\psi_{C}]$). Evaluation of the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ selection criteria is straightforward: We simply assess the conditional accuracies for the admitted points after stratifying by the predictions and the true labels, each in turn.
When Alg. 1 returns $q^{\prime}_{\rm{min}}=\infty$, we obtain a useful empirical indicator that the model is too weak, or the data insufficient, to reliably obtain class- and prediction-conditional estimates at the specified $\alpha$ value.
Algorithm 1 Search Algorithm to Find $q^{\prime}_{\rm{min}}$ to Determine the $\operatorname{\text{{high-reliability}}}$ Region
1:cached $(q^{\prime},\operatorname{\text{{sdm}}}({\bm{z}}^{\prime}))$ for $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, $\alpha\in(\frac{1}{C},1]$
2:procedure find-min-rescaled-similarity(cached $(q^{\prime},\operatorname{\text{{sdm}}}({\bm{z}}^{\prime}))$ for $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, $\alpha\in(\frac{1}{C},1]$)
3:  $q^{\prime}_{\rm{min}}\leftarrow\infty$ $\triangleright$ A finite $q^{\prime}_{\rm{min}}$ may not be found.
4:  $\psi_{1},\ldots\psi_{C}]\leftarrow[\infty,\ldots,\infty]$
5:  $\text{sortedList}~\leftarrow{\rm{sorted}}~[q^{\prime}\in\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}~{\rm{s.t.}}~\left\lfloor{q^{\prime}}\right\rfloor>0]$ $\triangleright$ Restricted to $\left\lfloor{q^{\prime}}\right\rfloor>0$ to exclude OOD
6:  for $\operatorname{\text{$q$}}^{*}\in\text{sortedList}$ do
7:    Construct ${\rm{eCDF}}^{y_{1}}_{\rm{ca}},\ldots,{\rm{eCDF}}^{y_{C}}_{\rm{ca}}$ for all $q^{\prime}\geq\operatorname{\text{$q$}}^{*}$ in $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\triangleright$ eCDFs for $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})$ (Eq. 6), stratified by $y$
8:    Calculate $\psi_{c}={\rm{inverseCDF}}^{y_{c}}_{\rm{ca}}(1-\alpha^{\prime})~\forall~c~\in\{1,\ldots,C\}$ $\triangleright$ Quantile functions are inverses of L. 7
9:    if ${\rm{all}}(~[\psi_{1},\ldots\psi_{C}]\geq\alpha~)$ then $\triangleright$ Element-wise comparison
10:     $q^{\prime}_{\rm{min}}\leftarrow\operatorname{\text{$q$}}^{*}$
11:     break
12:  return $q^{\prime}_{\rm{min}}$, $\psi_{1},\ldots\psi_{C}]$
[…]
4 Experiments
We provide controlled comparisons of our proposed methods over representative LLMs and tasks, systematically ablating relevant components, holding the data and underlying LLM constant, ceteris paribus. We consider in-distribution, co-variate shifted, and out-of-distribution test sets. We consider representative estimators over the existing LLM architecture (i.e., without additional parameters); with CNN adaptors; and with the $\operatorname{\text{{sdm}}}$ activation layer.
4.1 Task: Binary Sentiment Classification
$\operatorname{\text{{Sentiment}}}$: $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$.
Our first task is predicting the sentiment of movie reviews. We use the commonly used benchmark data of Maas et al. (2011). This is a binary classification task with $y\in\{0={\rm{negative}},1={\rm{positive}}\}$. $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ are constructed from a total of 18k instances. The held-out set for evaluation, $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=1583$, is from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. This is a well-studied task for which the surface-level signals correlated with the target labels are expected to be effectively modeled by large parameter LLMs; as such, relatively high task accuracies are expected. We use the label $\operatorname{\text{{Sentiment}}}$ for the in-distribution test set.
$\operatorname{\text{{SentimentOOD}}}$.
To evaluate the behavior of the estimators over out-of-distribution (OOD) data, we consider an additional evaluation set, $\operatorname{\text{{SentimentOOD}}}$, $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=4750$. We use the SemEval-2017 Task 4a test set (Rosenthal et al., 2017), which consists of short-form social media posts that differ in the distribution of topics, language styles, and lengths relative to the movie reviews. We balance the test set, dropping the third class (neutral), setting the semantics of the true labels to be the same as that of the movie reviews: $y\in\{0={\rm{negative}},1={\rm{positive}}\}$.
$\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$.
In the Appendix, we consider two additional out-of-distribution challenge test sets, $\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$, constructed by randomly shuffling the input documents for each of $\operatorname{\text{{Sentiment}}}$ and $\operatorname{\text{{SentimentOOD}}}$, respectively. The semantics of the original labels are maintained. This requires the models and estimators to attempt a sentiment classification over the bag-of-words input, or reject the classification. This represents the setting where an LLM is given far out-of-distribution input, and additionally provides a control on test-set contamination of the underlying LLMs, which due to the shuffling, are relatively unlikely to have seen all of the long, contiguous n-gram sequences from these documents in training or fine-tuning.
Prompt.
For this task, we prompt the LLMs for a binary classification (Yes or No) as follows: Here is a movie review. <review> DOCUMENT </review> Is the sentiment of the movie review positive? Answer Yes if the sentiment is positive. Answer No if the sentiment is negative. Start your response with Yes or No.
We replace DOCUMENT with the corresponding text for each instance.
4.2 Task: $\operatorname{\text{{Factcheck}}}$
$\operatorname{\text{{Factcheck}}}$.
As a more challenging binary classification task for LLMs, we consider the fact check data of Azaria & Mitchell (2023). The training and calibration sets, a combined total of 6k instances, consist of single sentence statements that have been semi-automatically generated via templates and a knowledge base. The task is to determine whether the statement is true or false, $y\in\{0={\rm{false}},1={\rm{true}}\}$. The held-out eval set ($\operatorname{\text{{Factcheck}}}$), $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=245$, the focus of our analysis, has been constructed by having an LLM generate a statement continued from a true statement not otherwise in the dataset. These evaluation statements are checked manually and assigned labels by human annotators. In addition to being a relatively challenging task that evaluates—at least in principle—the latent knowledge stored within an LLM’s parameters, the test set is representative of the types of co-variate shifts over high-dimensional inputs that can be problematic for real applications, and challenging to characterize without model assistance and ground-truth labels. It was observed in Azaria & Mitchell (2023) that the accuracy of existing LLM classifiers is lower on this generated, held-out test set compared to the calibration set. However, these test sentences would seem to also be simple true-false statements, reflecting that it is not always immediately obvious for a human user to detect distribution shifts over high-dimensional inputs. As such, we seek for our models and estimators to reflect such shifts via the predictive uncertainty.
$\operatorname{\text{{FactcheckShuffled}}}$.
As with the sentiment task, in the Appendix we also consider an additional out-of-distribution challenge test set, $\operatorname{\text{{FactcheckShuffled}}}$, constructed by randomly shuffling the input documents of $\operatorname{\text{{Factcheck}}}$. Since the task prompt (below) seeks to classify errors, we set the ground-truth labels of the shuffled counterparts to $y=0$.
Prompt.
Similar to the sentiment task, we prompt the LLMs for a binary classification (Yes or No) as follows: Here is a statement that may contain errors. <statement> DOCUMENT </statement> Is the statement true? Answer Yes if the statement is true. Answer No if the statement is false. Start your response with Yes or No.
We replace DOCUMENT with the corresponding text for each instance.
4.3 Models
We consider two representative, publicly-available decoder-only Transformer-based language models: The 3.8 billion-parameter Phi-3.5-mini-instruct model ($\operatorname{\text{{phi3.5}}}$) (Abdin et al., 2024), and the 47 billion-parameter Mixtral 8x7B Instruct v0.1 mixture-of-experts model ($\operatorname{\text{{Mixtral8x7B}}}$) (Jiang et al., 2024). The parameters of these models stay fixed in all experiments.
Hidden-states.
For both models, we take as $\bm{h}$} the concatenation of the final-layer hidden state of the final sequence position (i.e., the hidden state that is the input to the linear-layer over the output vocabulary for the Yes or No generation) with the sequence-wise mean over all final-layer hidden states. For $\operatorname{\text{{phi3.5}}}$, this results in ${\bm{h}}\in\mathbb{R}^{6144}$, and for $\operatorname{\text{{Mixtral8x7B}}}$, ${\bm{h}}\in\mathbb{R}^{8192}$.
4.4 Estimators
Holding each underlying LLM constant, we examine representative classes of estimators used with neural networks, setting $\alpha=0.95$ for all experiments. At the most basic, but also, perhaps the most commonly used in practice, representing the absence of a post-hoc calibration method, we simply threshold the output, $\operatorname{softmax}({\bm{z}})\geq\alpha$, where the temperature $\tau=1$. As an established empirical approach for calibrating neural networks, we provide a comparison to temperature scaling (Guo et al., 2017), a single parameter version of post-hoc Platt-scaling (Platt, 1999), with the label $\operatorname{\text{{tempScaling}}}$. In this case, the estimator is the thresholding of the output $\operatorname{softmax}({\bm{z}};\tau)\geq\alpha$ after learning a value for $\tau$ over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. We also provide a comparison to two representative conformal predictors, the $\operatorname{\textsc{APS}}$ method of Romano et al. (2020) and the adaptiveness-optimized $\operatorname{\textsc{RAPS}}$ algorithm of Angelopoulos et al. (2021). The admission criteria for the $\operatorname{\textsc{APS}}$ and $\operatorname{\textsc{RAPS}}$ estimators is prediction sets of size 1, at the $0.05$ level (i.e., $1-\alpha$, as defined here). We consider these estimators over the logits corresponding to the Yes and No indexes of the output linear-layer of the underlying LLM ($\operatorname{\text{{phi3.5}}}$ and $\operatorname{\text{{Mixtral8x7B}}}$), which provides a reference point without introducing additional adaptor layers. We also consider these baselines over 1-D CNN adaptors over $\bm{h}$} of each LLM ($\operatorname{\text{{phi3.5+adaptor}}}$ and $\operatorname{\text{{Mixtral8x7B+adaptor}}}$). These baseline adaptors are identical to those used in the corresponding $\operatorname{\text{{sdm}}}$ activation layers, with $M=1000$, and are similarly trained for $J=10$ iterations of 200 epochs, but unlike the $\operatorname{\text{{sdm}}}$ activation layers, the stopping criteria is the maximum balanced (across classes) average cross-entropy loss.
We then compare to the final-layer $\operatorname{\text{{sdm}}}$ activations over $\bm{h}$} of each LLM ($\operatorname{\text{{phi3.5+sdm}}}$ and $\operatorname{\text{{Mixtral8x7B+sdm}}}$). For reference, we provide the result of thresholding a $\operatorname{softmax}$ over these adaptors at $\alpha$, as above, as well as a thresholded $\operatorname{softmax}$ that simply treats $\operatorname{\text{$d$}}$ as the inverse-temperature, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$, which is equivalent to setting q=e2q=e-2 in the $\operatorname{\text{{sdm}}}$ activation. We also consider an analogous threshold over the activation output, $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})\geq\alpha$, for which we use $\operatorname{\text{{sdm}}}$ as the estimator label. Finally, we use the label $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ for the selection criteria proposed in Eq. 9.
As a common point of reference, the label $\operatorname{\text{{no-reject}}}$ refers to predictions without any selective filtering (i.e., the raw output accuracies, derived from the $\operatornamewithlimits{arg\,max}$ over the final linear-layer).
5 Results
In-distribution data.
The results for the in-distribution $\operatorname{\text{{Sentiment}}}$ dataset appear in Table 1. Even on this in-distribution dataset, the estimators over the underlying LLMs without adaptor layers exhibit modest over-confidence, which is reflected in conditional accuracies that fall below the expected $\alpha$. The estimators over the adaptor layers and the $\operatorname{\text{{sdm}}}$ activation all obtain the desired conditional accuracies, with the class-wise accuracies of the models themselves $\geq\alpha$, with differences arising in the proportion of admitted points. Here and elsewhere, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ tends to be overly conservative in rejecting points. To be expected, the $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ estimator tends to be more conservative than simply thresholding the $\operatorname{\text{{sdm}}}$ activation on the in-distribution data, but the latter lacks the assurances on the recall obtained by the constraints on the $\operatorname{\text{{high-reliability}}}$ region. In practice, this behavior can be used as a basis to triage selective classifications: For example, documents in the $\operatorname{\text{{high-reliability}}}$ region might be treated as automated, or semi-automated, predictions in the decision pipeline, whereas other documents might be triaged by $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})_{\hat{y}}$ for calling more resource-intensive LLM tools, or human adjudication. Importantly, as we discuss below with the co-variate-shifted and out-of-distribution datasets, just using the $\operatorname{softmax}$ or the other estimators does not provide a reliable substrate for basing such conditional branching decisions.
Table 1: Comparison of estimators for the sentiment datasets, with $\alpha$=0.95=0.95. R indicates all predictions were rejected, which is preferred over falling under the expected accuracy. $n=|\text{Admitted}|$, the count of non-rejected documents.
Class-conditional Prediction-conditional Marginal
$y=0$ $y=1$ $\hat{y}=0$ $\hat{y}=1$ $y\in\{0,1\}$
Dataset Model Estimator Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{no-reject}}}$ 0.98 0.50 0.85 0.50 0.86 0.57 0.98 0.43 0.91 1.
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{softmax}}}$ 0.98 0.50 0.86 0.48 0.88 0.56 0.98 0.42 0.93 0.98
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{tempScaling}}}$ 0.99 0.49 0.91 0.41 0.93 0.52 0.99 0.38 0.95 0.90
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{APS}}$ 0.99 0.49 0.92 0.40 0.94 0.51 0.99 0.37 0.96 0.89
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{RAPS}}$ 0.99 0.48 0.91 0.41 0.93 0.51 0.99 0.38 0.95 0.90
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.97 0.50 0.95 0.50 0.96 0.51 0.97 0.49 0.96 1.
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.99 0.42 1.00 0.42 1.00 0.42 0.99 0.42 0.99 0.84
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.99 0.42 1.00 0.41 1.00 0.42 0.99 0.41 0.99 0.83
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.98 0.45 0.98 0.45 0.98 0.45 0.98 0.45 0.98 0.90
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.98 0.45 0.98 0.44 0.98 0.45 0.98 0.44 0.98 0.89
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.96 0.50 0.96 0.50 0.96 0.50 0.96 0.50 0.96 1.
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.97 0.48 0.97 0.48 0.97 0.48 0.97 0.48 0.97 0.96
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ 0.99 0.30 0.99 0.24 1.00 0.30 0.99 0.24 0.99 0.54
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.99 0.43 0.99 0.38 0.99 0.43 0.99 0.38 0.99 0.81
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ 1.00 0.37 0.99 0.30 0.99 0.38 1.00 0.30 0.99 0.68
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{no-reject}}}$ 0.98 0.50 0.88 0.50 0.89 0.55 0.98 0.45 0.93 1.
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{softmax}}}$ 0.98 0.50 0.88 0.50 0.89 0.55 0.98 0.45 0.93 1.00
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{tempScaling}}}$ 0.99 0.50 0.90 0.48 0.91 0.54 0.98 0.44 0.94 0.98
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{APS}}$ 0.98 0.49 0.91 0.47 0.92 0.52 0.98 0.44 0.95 0.96
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{RAPS}}$ 0.99 0.49 0.92 0.47 0.93 0.52 0.98 0.44 0.95 0.96
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.97 0.50 0.96 0.50 0.96 0.51 0.97 0.49 0.97 1.
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.99 0.45 0.99 0.43 0.99 0.45 0.99 0.43 0.99 0.87
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.99 0.43 0.99 0.41 0.99 0.43 0.99 0.41 0.99 0.84
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.99 0.46 0.98 0.45 0.98 0.46 0.99 0.44 0.99 0.91
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.99 0.46 0.98 0.45 0.98 0.47 0.98 0.45 0.98 0.92
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.96 0.50 0.95 0.50 0.95 0.51 0.96 0.49 0.96 1.
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.97 0.49 0.96 0.49 0.96 0.50 0.97 0.49 0.97 0.98
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ 0.99 0.43 0.99 0.33 0.99 0.43 0.99 0.33 0.99 0.77
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.98 0.48 0.98 0.43 0.98 0.47 0.98 0.43 0.98 0.90
$\operatorname{\text{{Sentiment}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ 0.99 0.41 0.98 0.33 0.99 0.41 0.98 0.33 0.99 0.74
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{no-reject}}}$ 1.00 0.50 0.53 0.50 0.68 0.73 0.99 0.27 0.76 1.
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{softmax}}}$ 1.00 0.50 0.54 0.46 0.70 0.71 0.99 0.25 0.78 0.96
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{tempScaling}}}$ 1.00 0.49 0.58 0.30 0.80 0.62 0.99 0.17 0.84 0.79
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{APS}}$ 1.00 0.49 0.59 0.28 0.81 0.60 0.99 0.17 0.85 0.77
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{RAPS}}$ 1.00 0.49 0.59 0.28 0.81 0.60 0.99 0.17 0.85 0.77
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.47 0.50 0.70 0.50 0.61 0.38 0.57 0.62 0.59 1.
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.57 0.03 0.96 0.07 0.84 0.02 0.85 0.07 0.85 0.09
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.60 0.02 0.97 0.05 0.86 0.01 0.87 0.06 0.87 0.07
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.46 0.14 0.83 0.18 0.67 0.09 0.68 0.22 0.68 0.32
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.48 0.13 0.82 0.18 0.66 0.10 0.68 0.22 0.68 0.32
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.92 0.50 0.84 0.50 0.85 0.54 0.91 0.46 0.88 1.
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.96 0.42 0.87 0.45 0.87 0.46 0.96 0.41 0.91 0.87
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ 1. <0.01 1. <0.01 1. <0.01 1. <0.01 1. <0.01
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. 0.01 0.98 0.01 0.98 0.01 1. 0.01 0.99 0.02
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ 1. <0.01 1. <0.01 1. <0.01 1. <0.01 1. 0.01
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{no-reject}}}$ 1.00 0.50 0.35 0.50 0.61 0.82 1.00 0.18 0.67 1.
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{softmax}}}$ 1.00 0.50 0.35 0.49 0.61 0.82 1.00 0.17 0.68 0.99
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{tempScaling}}}$ 1.00 0.49 0.37 0.41 0.66 0.75 0.99 0.15 0.71 0.90
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{APS}}$ 1.00 0.45 0.44 0.32 0.71 0.63 0.99 0.14 0.77 0.77
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{RAPS}}$ 1.00 0.45 0.44 0.32 0.72 0.63 0.99 0.14 0.77 0.77
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.88 0.50 0.51 0.50 0.64 0.69 0.82 0.31 0.70 1.
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.98 0.02 0.83 0.07 0.66 0.04 0.99 0.06 0.87 0.10
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.98 0.01 0.90 0.05 0.67 0.02 1.00 0.05 0.91 0.06
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.94 0.14 0.63 0.18 0.67 0.20 0.93 0.12 0.77 0.32
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.94 0.14 0.63 0.18 0.67 0.20 0.93 0.12 0.76 0.32
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.71 0.50 0.83 0.50 0.81 0.44 0.74 0.56 0.77 1.
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.74 0.43 0.86 0.47 0.83 0.39 0.78 0.52 0.80 0.91
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ 1. <0.01 0.98 0.02 0.78 <0.01 1. 0.02 0.98 0.02
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.98 0.05 0.96 0.04 0.97 0.05 0.98 0.04 0.97 0.08
$\operatorname{\text{{SentimentOOD}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ 0.9487 0.01 0.96 0.01 0.9487 0.01 0.96 0.01 0.95 0.02
Co-variate-shifted and Out-of-distribution data.
The bottom half of Table 1 provides the results for $\operatorname{\text{{SentimentOOD}}}$. Here, the distinctions between the estimators become clear, with the non-$\operatorname{\text{{sdm}}}$-based estimators performing poorly, even in terms of the marginal accuracy. That would come as a surprise to end-users, whereas with the $\operatorname{\text{{sdm}}}$ and $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ estimators, the out-of-distribution documents are more reliably rejected, with the few admitted predictions generally obtaining high conditional accuracies, despite the relatively low accuracies of the test set without selection (see $\operatorname{\text{{no-reject}}}$). A similar pattern is observed over the $\operatorname{\text{{Factcheck}}}$ dataset in Table 2.
Understanding $q^{\prime}_{\rm{min}}$.
For reference, Table 3 provides the results over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for the $\operatorname{\text{{sdm}}}$-based estimates. The value of $q^{\prime}_{\rm{min}}$ tends to increase as the accuracy over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ decreases, reflecting a more conservative $\operatorname{\text{{high-reliability}}}$ region that admits fewer points. Alg. 1 failed to find a finite $q^{\prime}_{\rm{min}}$ for $\operatorname{\text{{Mixtral8x7B+sdm}}}$ over the $\operatorname{\text{{Factcheck}}}$ calibration set at $\alpha=0.95$, so for reference, we also show the $\operatorname{\text{{high-reliability}}}$ region at $\alpha=0.94$, as well as in Table 2. In this way, $q^{\prime}_{\rm{min}}$ provides a principled, data-driven indicator of the reliability of the estimates, which is interpretable as a simple indicator as to whether the conditional accuracies are, or are not, obtainable over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ at the specified $\alpha$.
Far out-of-distribution data.
The Appendix provides Table 4 and Table 5 with analogous results for the $\operatorname{\text{{SentimentShuffled}}}$, $\operatorname{\text{{SentimentOODShuffled}}}$, and $\operatorname{\text{{FactcheckShuffled}}}$ datasets. The selection criteria of Eq. 9 reliably rejects the challenging predictions, whereas the non-$\operatorname{\text{{sdm}}}$-based estimators fare poorly, in general. In this way, the $\operatorname{\text{{sdm}}}$ activation serves as an effective out-of-distribution detection method. With existing methods, defining an out-of-distribution point has been task- and problem-specific, and generally challenging over high-dimensional inputs, typically requiring additional modeling beyond that of the calibration or selection method. In contrast, Eq. 9 provides a principled approach for determining such cut-offs in a data- and model-driven manner, with minimal hyper-parameters, resulting in a separation of points over which the estimator tends to be reliable (namely, the admitted points) and those over which the estimates themselves tend to be unreliable.
Table 2: Comparison of estimators for the factcheck datasets. Unless specified otherwise, $\alpha$=0.95=0.95. R indicates all predictions were rejected, which is preferred over falling under the expected accuracy. $n=|\text{Admitted}|$, the count of non-rejected documents.
Class-conditional Prediction-conditional Marginal
$y=0$ $y=1$ $\hat{y}=0$ $\hat{y}=1$ $y\in\{0,1\}$
Dataset Model Estimator Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{no-reject}}}$ 0.94 0.51 0.71 0.49 0.78 0.62 0.92 0.38 0.83 1.
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{softmax}}}$ 0.94 0.51 0.73 0.46 0.79 0.60 0.92 0.36 0.84 0.97
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{tempScaling}}}$ 0.97 0.38 0.79 0.37 0.83 0.45 0.96 0.31 0.88 0.76
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{APS}}$ 0.98 0.22 0.82 0.27 0.82 0.27 0.98 0.23 0.89 0.50
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{RAPS}}$ 0.98 0.20 0.84 0.28 0.81 0.24 0.98 0.24 0.90 0.47
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.33 0.51 0.94 0.49 0.85 0.20 0.57 0.80 0.62 1.
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.40 0.08 0.99 0.33 0.89 0.04 0.87 0.37 0.87 0.41
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.38 0.07 0.99 0.29 0.86 0.03 0.88 0.33 0.88 0.36
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.26 0.14 0.99 0.38 0.90 0.04 0.78 0.48 0.79 0.52
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.36 0.18 0.98 0.35 0.89 0.07 0.74 0.46 0.76 0.53
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.70 0.51 0.88 0.49 0.86 0.42 0.73 0.58 0.79 1.
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.75 0.27 0.94 0.39 0.89 0.22 0. 85 0.43 0.86 0.65
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ R 0. 1. 0.03 R 0. 1. 0.03 1. 0.03
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. 0.01 0.97 0.14 0.75 0.02 1. 0.14 0.97 0.16
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ R 0. 1. 0.12 R 0. 1. 0.12 1. 0.12
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{no-reject}}}$ 0.98 0.51 0.48 0.49 0.66 0.76 0.95 0.24 0.73 1.
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{softmax}}}$ 0.98 0.51 0.48 0.49 0.66 0.76 0.95 0.24 0.73 1.
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{tempScaling}}}$ 0.99 0.50 0.46 0.43 0.68 0.73 0.98 0.20 0.75 0.93
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{APS}}$ 1. 0.18 0.80 0.16 0.84 0.21 1. 0.13 0.90 0.34
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{RAPS}}$ 1. 0.14 0.66 0.20 0.67 0.21 1. 0.13 0.80 0.35
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.56 0.51 0.87 0.49 0.82 0.36 0.65 0.64 0.71 1.
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.68 0.11 0.97 0.31 0.90 0.09 0.89 0.34 0.89 0.42
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.70 0.09 0.97 0.29 0.89 0.07 0.91 0.31 0.91 0.39
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.62 0.22 0.96 0.37 0.89 0.16 0.80 0.44 0.83 0.59
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.65 0.22 0.96 0.35 0.92 0.16 0.81 0.41 0.84 0.57
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.63 0.51 0.90 0.49 0.87 0.38 0.70 0.62 0.76 1.
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.67 0.34 0.96 0.40 0.93 0.24 0.78 0.50 0.83 0.74
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ R 0. 0.80 0.04 0. 0.01 1. 0.03 0.80 0.04
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.88 0.10 0.95 0.18 0.91 0.09 0.93 0.18 0.93 0.27
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ R 0. R 0. R 0. R 0. R 0.
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}},\alpha=0.94$ 0.85 0.11 0.95 0.18 0.92 0.10 0.91 0.19 0.91 0.29
$\operatorname{\text{{Factcheck}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}},\alpha=0.94$ 1. 0.03 0.95 0.16 0.80 0.04 1. 0.15 0.96 0.19
Table 3: Reference results over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ to illustrate the behavior of $q^{\prime}_{\rm{min}}$. The value of $q^{\prime}_{\rm{min}}$ tends to increase as the accuracy over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ decreases, reflecting a more conservative $\operatorname{\text{{high-reliability}}}$ region. Alg. 1 failed to find a finite $q^{\prime}_{\rm{min}}$ for $\operatorname{\text{{Mixtral8x7B+sdm}}}$ over the $\operatorname{\text{{Factcheck}}}$ calibration set at $\alpha=0.95$, so for reference, we also show the $\operatorname{\text{{high-reliability}}}$ region at $\alpha=0.94$.
Class-conditional Prediction-conditional Marginal
$y=0$ $y=1$ $\hat{y}=0$ $\hat{y}=1$ $y\in\{0,1\}$
Dataset Model Estimator Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.95 0.50 0.96 0.50 0.96 0.50 0.95 0.50 0.96 1.
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.96 0.48 0.97 0.48 0.97 0.47 0.96 0.49 0.97 0.96
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ 0.99 0.31 0.99 0.24 1.00 0.31 0.99 0.24 0.99 0.55
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.99 0.42 0.99 0.39 0.99 0.42 0.99 0.39 0.99 0.81
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}},\alpha=0.95,{\color[rgb]{0.2,0.2,1}q^{\prime}_{\rm{min}}=52.2}$ 0.99 0.38 0.99 0.32 0.99 0.38 0.99 0.31 0.99 0.69
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.96 0.50 0.96 0.50 0.96 0.50 0.96 0.50 0.96 1.
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.96 0.49 0.97 0.49 0.97 0.49 0.96 0.49 0.97 0.98
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ 1.00 0.43 0.99 0.35 0.99 0.43 1.00 0.34 0.99 0.78
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.99 0.47 0.98 0.43 0.98 0.47 0.98 0.43 0.98 0.90
$\operatorname{\text{{Sentiment}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}},\alpha=0.95,{\color[rgb]{0.2,0.2,1}q^{\prime}_{\rm{min}}=63.0}$ 0.99 0.41 0.99 0.34 0.99 0.41 0.99 0.34 0.99 0.74
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.90 0.50 0.91 0.50 0.91 0.49 0.90 0.51 0.90 1.
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.94 0.32 0.96 0.41 0.95 0.31 0.96 0.41 0.96 0.72
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ 0.98 0.08 1.00 0.07 1.00 0.08 0.98 0.07 0.99 0.15
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.98 0.33 0.99 0.27 0.99 0.33 0.98 0.27 0.98 0.60
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}},\alpha=0.95,{\color[rgb]{0.2,0.2,1}q^{\prime}_{\rm{min}}=95.0}$ 1.00 0.19 0.99 0.12 1.00 0.19 0.99 0.12 1.00 0.31
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.90 0.50 0.92 0.50 0.92 0.49 0.90 0.51 0.91 1.
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.94 0.44 0.96 0.45 0.96 0.43 0.94 0.46 0.95 0.89
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ 0.98 0.28 0.98 0.17 0.99 0.27 0.96 0.17 0.98 0.44
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 0.97 0.41 0.95 0.32 0.96 0.41 0.95 0.32 0.96 0.73
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}},\alpha=0.95,{\color[rgb]{0.2,0.2,1}q^{\prime}_{\rm{min}}=\infty}$ R 0. R 0. R 0. R 0. R 0.
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}},\alpha=0.94$ 0.96 0.42 0.95 0.33 0.96 0.42 0.95 0.32 0.96 0.74
$\operatorname{\text{{Factcheck}}}\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}},\alpha=0.94,{\color[rgb]{0.2,0.2,1}q^{\prime}_{\rm{min}}=134.0}$ 1.00 0.33 0.96 0.14 0.99 0.33 0.99 0.13 0.99 0.46
6 Conclusion
We introduced $\operatorname{\text{{sdm}}}$ activation functions, which provide more robust estimators of the predictive uncertainty than the commonly-used $\operatorname{softmax}$ function. In this way, $\operatorname{\text{{sdm}}}$ activations provide a principled, data-driven substrate for approaching selective classification, calibration, and out-of-distribution detection with LLMs.
Appendix A Appendix
We provide the results for the far out-of-distribution (OOD) shuffled datasets in Section A.1. Additional implementation details are included in Section A.2.
A.1 Far OOD Shuffled Datasets
Table 4 shows results for the $\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$ datasets, and Table 5 shows results for the $\operatorname{\text{{FactcheckShuffled}}}$ datasets, as discussed in Section 5.
Table 4: Comparison of estimators for the shuffled sentiment datasets, with $\alpha$=0.95=0.95. R indicates all predictions were rejected, which is preferred over falling under the expected accuracy. $n=|\text{Admitted}|$, the count of non-rejected documents.
Class-conditional Prediction-conditional Marginal
$y=0$ $y=1$ $\hat{y}=0$ $\hat{y}=1$ $y\in\{0,1\}$
Dataset Model Estimator Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{no-reject}}}$ 1.00 0.50 0.18 0.50 0.55 0.91 0.98 0.09 0.59 1.
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{softmax}}}$ 1.00 0.50 0.16 0.45 0.57 0.88 0.99 0.07 0.60 0.95
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{tempScaling}}}$ 1. 0.44 0.12 0.18 0.74 0.60 1. 0.02 0.75 0.62
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{APS}}$ 1.00 0.42 0.14 0.15 0.76 0.55 0.97 0.02 0.77 0.57
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{RAPS}}$ 1. 0.41 0.13 0.16 0.75 0.55 1. 0.02 0.76 0.57
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.83 0.50 0.81 0.50 0.81 0.51 0.82 0.49 0.82 1.
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.98 0.13 0.97 0.12 0.98 0.13 0.97 0.12 0.97 0.25
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.98 0.11 0.97 0.10 0.97 0.11 0.97 0.10 0.97 0.21
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.93 0.23 0.90 0.24 0.90 0.24 0.93 0.23 0.91 0.48
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.92 0.24 0.91 0.24 0.91 0.24 0.92 0.24 0.92 0.48
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.99 0.50 0.32 0.50 0.59 0.83 0.97 0.17 0.66 1.
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.99 0.49 0.29 0.36 0.65 0.74 0.98 0.11 0.69 0.85
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ R 0. R 0. R 0. R 0. R 0.
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. 0.01 1. <0.01 1. 0.01 1. <0.01 1. 0.01
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ 1. <0.01 1. <0.01 1. <0.01 1. <0.01 1. <0.01
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{no-reject}}}$ 0.99 0.50 0.35 0.50 0.60 0.82 0.97 0.18 0.67 1.
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{softmax}}}$ 0.99 0.50 0.34 0.48 0.61 0.81 0.97 0.17 0.67 0.98
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{tempScaling}}}$ 0.99 0.48 0.37 0.38 0.67 0.72 0.98 0.14 0.72 0.86
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{APS}}$ 0.99 0.45 0.44 0.27 0.75 0.60 0.97 0.12 0.79 0.72
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{RAPS}}$ 0.99 0.44 0.43 0.29 0.73 0.60 0.98 0.13 0.77 0.73
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.91 0.50 0.72 0.50 0.76 0.60 0.89 0.40 0.81 1.
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.99 0.28 0.83 0.14 0.92 0.31 0.98 0.12 0.94 0.43
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.99 0.26 0.82 0.11 0.93 0.28 0.98 0.09 0.94 0.37
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.97 0.34 0.78 0.24 0.86 0.38 0.95 0.19 0.89 0.58
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.96 0.35 0.79 0.23 0.87 0.38 0.93 0.20 0.89 0.58
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.79 0.50 0.82 0.50 0.82 0.48 0.79 0.52 0.80 1.
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.80 0.48 0.83 0.49 0.82 0.47 0.80 0.50 0.81 0.97
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ R 0. R 0. R 0. R 0. R 0.
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. <0.01 R 0. 1. <0.01 R 0. 1. <0.01
$\operatorname{\text{{SentimentShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ R 0. R 0. R 0. R 0. R 0.
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{no-reject}}}$ 1.00 0.50 0.35 0.50 0.60 0.82 0.99 0.18 0.67 1.
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{softmax}}}$ 1.00 0.50 0.34 0.47 0.62 0.81 0.99 0.16 0.68 0.97
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{tempScaling}}}$ 1.00 0.49 0.34 0.29 0.72 0.68 0.99 0.10 0.75 0.78
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{APS}}$ 1.00 0.48 0.36 0.27 0.74 0.65 0.99 0.10 0.77 0.75
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{RAPS}}$ 1.00 0.49 0.36 0.27 0.74 0.66 0.99 0.10 0.77 0.76
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.64 0.50 0.64 0.50 0.64 0.50 0.64 0.50 0.64 1.
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.78 0.01 0.93 0.03 0.81 0.01 0.92 0.03 0.89 0.04
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.79 0.01 0.94 0.03 0.83 0.01 0.93 0.03 0.90 0.03
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.68 0.11 0.73 0.14 0.67 0.12 0.74 0.14 0.71 0.26
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.70 0.12 0.72 0.14 0.68 0.12 0.74 0.13 0.71 0.25
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.97 0.50 0.63 0.50 0.72 0.67 0.95 0.33 0.80 1.
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.98 0.47 0.64 0.43 0.75 0.62 0.97 0.28 0.82 0.89
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ R 0. R 0. R 0. R 0. R 0.
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. <0.01 1. <0.01 1. <0.01 1. <0.01 1. 0.01
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ 1. <0.01 1. <0.01 1. <0.01 1. <0.01 1. <0.01
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{no-reject}}}$ 1.00 0.50 0.12 0.50 0.53 0.94 1.00 0.06 0.56 1.
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{softmax}}}$ 1.00 0.50 0.12 0.49 0.54 0.93 1.00 0.06 0.56 0.99
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{tempScaling}}}$ 1. 0.49 0.14 0.34 0.63 0.78 1. 0.05 0.65 0.83
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{APS}}$ 1. 0.36 0.24 0.18 0.72 0.51 1. 0.04 0.74 0.55
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{RAPS}}$ 1. 0.36 0.23 0.19 0.72 0.51 1. 0.04 0.74 0.55
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.97 0.50 0.25 0.50 0.56 0.86 0.89 0.14 0.61 1.
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 1. 0.04 0.23 0.03 0.62 0.07 1. 0.01 0.66 0.07
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 1. 0.02 0.27 0.02 0.58 0.03 1. 0.01 0.64 0.04
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.99 0.18 0.21 0.15 0.60 0.29 0.94 0.03 0.64 0.32
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.99 0.18 0.21 0.14 0.61 0.29 0.93 0.03 0.64 0.32
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.75 0.50 0.71 0.50 0.72 0.52 0.74 0.48 0.73 1.
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.79 0.43 0.73 0.45 0.73 0.46 0.78 0.42 0.76 0.89
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ 1. <0.01 1. <0.01 1. <0.01 1. <0.01 1. <0.01
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. <0.01 0.83 <0.01 0.88 0.01 1. <0.01 0.93 0.01
$\operatorname{\text{{SentimentOODShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ 1. <0.01 1. <0.01 1. <0.01 1. <0.01 1. <0.01
Table 5: Comparison of estimators for the shuffled factcheck datasets. Unless specified otherwise, $\alpha$=0.95=0.95. R indicates all predictions were rejected, which is preferred over falling under the expected accuracy. $n=|\text{Admitted}|$, the count of non-rejected documents.
Class-conditional Prediction-conditional Marginal
$y=0$ $y=1$ $\hat{y}=0$ $\hat{y}=1$ $y\in\{0,1\}$
Dataset Model Estimator Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$ Acc. $\frac{n}{|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|}$
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{no-reject}}}$ 0.91 1. - 0. 1. 0.91 0. 0.09 0.91 1.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{softmax}}}$ 0.92 0.99 - 0. 1. 0.91 0. 0.08 0.92 0.99
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\text{{tempScaling}}}$ 0.93 0.87 - 0. 1. 0.81 0. 0.06 0.93 0.87
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{APS}}$ 0.93 0.45 - 0. 1. 0.42 0. 0.03 0.93 0.45
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5}}}$ $\operatorname{\textsc{RAPS}}$ 0.95 0.52 - 0. 1. 0.50 0. 0.02 0.95 0.52
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.34 1. - 0. 1. 0.34 0. 0.66 0.34 1.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.20 0.24 - 0. 1. 0.05 0. 0.19 0.20 0.24
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.13 0.19 - 0. 1. 0.02 0. 0.17 0.13 0.19
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.24 0.38 - 0. 1. 0.09 0. 0.29 0.24 0.38
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.27 0.39 - 0. 1. 0.11 0. 0.29 0.27 0.39
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.66 1. - 0. 1. 0.66 0. 0.34 0.66 1.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.69 0.64 - 0. 1. 0.44 0. 0.20 0.69 0.64
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ R 0. - 0. R 0. R 0. R 0.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}$ R 0. - 0. R 0. R 0. R 0.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{phi3.5+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ R 0. - 0. R 0. R 0. R 0.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{no-reject}}}$ 0.98 1. - 0. 1. 0.98 0. 0.02 0.98 1.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{softmax}}}$ 0.98 1. - 0. 1. 0.98 0. 0.02 0.98 1.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\text{{tempScaling}}}$ 0.98 0.98 - 0. 1. 0.96 0. 0.02 0.98 0.98
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{APS}}$ 0.98 0.18 - 0. 1. 0.18 0. <0.01 0.98 0.18
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B}}}$ $\operatorname{\textsc{RAPS}}$ 0.98 0.23 - 0. 1. 0.23 0. <0.01 0.98 0.23
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{no-reject}}}$ 0.79 1. - 0. 1. 0.79 0. 0.21 0.79 1.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{softmax}}}$ 0.69 0.13 - 0. 1. 0.09 0. 0.04 0.69 0.13
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\text{{tempScaling}}}$ 0.55 0.09 - 0. 1. 0.05 0. 0.04 0.55 0.09
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{APS}}$ 0.77 0.40 - 0. 1. 0.31 0. 0.09 0.77 0.40
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+adaptor}}}$ $\operatorname{\textsc{RAPS}}$ 0.79 0.39 - 0. 1. 0.31 0. 0.08 0.79 0.39
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{no-reject}}}$ 0.76 1. - 0. 1. 0.76 0. 0.24 0.76 1.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}$ 0.79 0.65 - 0. 1. 0.51 0. 0.13 0.79 0.65
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$ R 0. - 0. R 0. R 0. R 0.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}$ 1. 0.01 - 0. 1. 0.01 R 0. 1. 0.01
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ R 0. - 0. R 0. R 0. R 0.
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}},\alpha=0.94$ 1. 0.01 - 0. 1. 0.01 R 0. 1. 0.01
$\operatorname{\text{{FactcheckShuffled}}}$ $\operatorname{\text{{Mixtral8x7B+sdm}}}$ $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}},\alpha=0.94$ 1. 0.01 - 0. 1. 0.01 R 0. 1. 0.01
A.2 Additional Implementation Details
The code, data, and models to replicate and reproduce our results are available at the following URL: https://github.com/ReexpressAI/sdm_activations
We mean center the input to $g$, the 1-D CNN of the $\operatorname{\text{{sdm}}}$ activation layer and the otherwise identical CNN adaptors of the baseline comparison estimators, via the mean and standard deviation over $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$. In all experiments with adaptor layers, $M=1000$ and we use a mini-batch size of 50. We use the Adam optimizer (Kingma & Ba, 2017) with a learning rate of $1\times 10^{-5}$ for training.
A.2.1 Implementation of the SDM Activation Function
As is typical with implementations of the $\operatorname{softmax}$ function, for numerical stability, rather than directly calculating $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})$, we instead use the mathematically equivalent $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime}-\max({\bm{z}}^{\prime}))$, shifting the input vector by its maximum value.
A.2.2 Implementation of the Empirical CDF Function
The conventions for implementing the empirical CDF functions follow in the expected ways, but we briefly highlight the key considerations below, as they can impact the behavior of the estimators.
  1. 1.
    The distance quantiles should be exclusionary at the boundaries. When $\operatorname{\text{$d_{\rm{nearest}}$}}=0$, the $1-{\rm{eCDF}}^{\cdot}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})$ quantile should be 1, and when $\operatorname{\text{$d_{\rm{nearest}}$}}$ is greater than the maximum observed distance (across $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}$ and ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$, and across $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ for ${\bm{x}}\in\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$, the latter case only occurring during training), the $1-{\rm{eCDF}}^{\cdot}_{\rm{ca}}(\operatorname{\text{$d_{\rm{nearest}}$}})$ quantile should be 0.
  2. 2.
    For the quantiles over an $\operatorname{\text{{sdm}}}$ activation, as needed for Alg. 1, saturated values at the high-end should be assigned a quantile of 1.